OpenStudy (anonymous):

Kandice notices that 2/7 of a gallon of milk is on the top shelf of the refrigerator, and 3/5 of gallon of milk is on the bottom shelf. How much milk is in the refrigerator in all?

OpenStudy (anonymous):

make the denominators the same

OpenStudy (anonymous):

so lets make the denominator 5*7=35, right?

OpenStudy (anonymous):

to do that, you need to multiply both the numerator and denominator of 2/7 by 5, remember what you do to the top you must do to the bottom

OpenStudy (anonymous):

so you have 10/35

OpenStudy (anonymous):

then you need to multiply both in 3/5 by 7, so you have 21/35

OpenStudy (anonymous):

you following?

OpenStudy (anonymous):

yes a little

OpenStudy (anonymous):

k do you understand what ive done so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

to add them, the bottom number of both fractions have to be common, and the easiest way to find the common denominator is just to multiply them by each other

OpenStudy (anonymous):

so now that youve found the common denominator and multiplied both fractions by the required number, just add the fractions together

OpenStudy (anonymous):

\[\frac{ 10 }{ 35 }+\frac{ 21 }{ 35 }\]

OpenStudy (anonymous):

31/35 of a gallon at this point you would simplify if you could, but you cant simplify this

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

what is 3 1/4 + 4 1/3

OpenStudy (anonymous):

np :) any other questions?